OpenStudy (anonymous):

how much of a 20% acid solution must be added to 80ml of a 15% acid solution to obtain an 18% acid solution?

OpenStudy (anonymous):

80ml of a 15% acid solution has 80*15%=12ml of acid. You can write the equation as (12+.2x)/(80+x)=.18 That means that 12+.2x=.18(80+x), which means 12+.2x=14.4+.18x, so .02x=2.4, x=2.4/.02, x=120. So if we add 120ml of a 20% solution to 80ml of a 15% solution, we'll end up with a 200ml solution that has .2*120+.15*80=24+12=36ml of acid, and 36/200=.18, so we have our 18% solution.

OpenStudy (anonymous):

Thanks buddy